# Go-with-the-Flow: Motion-Controllable Video Diffusion Models Using Real-Time Warped Noise

### v3

@@ 4.4.1 Local object motion control
Baselines. We evaluate our video generation model against three state-of-the-art baselines, SG-I2V [37], MotionClone [33], and DragAnything [58], to benchmark its ability to accurately control object and camera movements derived from a given input template. One of the most recent works, SG-I2V, is an I2V model for object motion transfer guided by bounding box trajectories. We adapt our user-defined polygons to bounding boxes as its input.
Qualitative results. From  Fig. 5, Fig. 6, and our supplementary webpage, we observe:
@@
(2) We found that SG-I2V frequently misinterprets object-specific movements as global camera shifts, resulting in scene-wide translations rather than accurate object manipulations.
@@
(5) Our model consistently outperforms these baselines by maintaining high object fidelity and 3D consistency, even in scenarios with intricate or overlapping motions. Notably, our approach preserves object integrity and introduces plausible physical interactions, such as generating realistic splashes when moving a duck within a tub. Extensive user studies and quantitative evaluations validate our superior performance in motion consistency, visual fidelity, and overall realism.
@@ 4.4.2 Motion transfer and camera movement control
Baselines. For the motion transfer T2V scenario, we compare with the recent state-of-the-art methods Diffusion Motion Transfer (DMT) [63] and MotionClone [33]. For the motion transfer I2V, we compare only with MotionClone as DMT does not take an image as input.
@@
Table 2: Quantitative comparisons of motion transfer. $\Uparrow$/$\Downarrow$ indicates a higher/lower value is better. Bold indicates the best results.
VBench $\Uparrow$
FID CoTracker Pixel CLIP- CLIP- FVD Subject Background Motion Temporal
$\Downarrow$ mIoU $\Uparrow$ MSE $\Downarrow$ text $\Uparrow$ image $\Uparrow$ $\Downarrow$ consistency consistency smoothness flickering
Motion transfer T2V on DAVIS
DMT - 0.85 - 0.31 0.95 - 0.86 0.92 0.94 0.91
MotionClone - 0.75 - 0.32 0.93 - 0.78 0.89 0.86 0.81
Ours - 0.70 - 0.33 0.98 - 0.88 0.93 0.97 0.89
Motion transfer I2V on DAVIS
MotionClone 99.41 0.72 0.068 0.31 0.94 1839.13 0.75 0.85 0.92 0.87
Ours 78.62 0.74 0.053 0.31 0.97 1211.37 0.88 0.92 0.98 0.93
Ours (prompt only) 76.62 0.52 0.088 0.31 0.96 1357.46 0.85 0.91 0.96 0.92
Camera movement transfer I2V on DL3DV
MotionClone 82.67 0.71 0.10 0.33 0.94 $1$.92 0.74 0.85 0.91 0.86
Ours 48.41 0.83 0.046 0.32 0.97 340.06 0.88 0.92 0.97 0.93
Ours (lower LoRA rank) 110.42 0.79 0.061 0.31 0.96 1200.35 0.84 0.91 0.96 0.93
Ours (w/o noise degradation) 55.16 0.84 0.047 0.32 0.96 453.24 0.86 0.91 0.96 0.93
Camera movement transfer I2V on WonderJourney
MotionClone 177.87 0.81 0.10 0.32 0.96 1928.12 0.75 0.87 0.93 0.87
Ours 128.27 0.85 0.07 0.31 0.98 1548.53 0.82 0.91 0.98 0.94
Results. We present both qualitative and quantitative comparisons with baselines in Tab. 2, Fig. 7, Fig. 8, Fig. 9, Fig. 10, and our supplementary webpage. We observe:
@@
(2) Our superior camera movement control: On the DL3DV and WonderJourney datasets, which involve substantial camera movement, our method notably outperforms MotionClone in both motion fidelity and general video quality. This highlights our method’s ability to effectively replicate intricate camera movements while maintaining visual coherence. For our depth-warping example Fig. 9, our results are far better than simply warping an image from its depth map, resulting in a smooth, realistic camera trajectory. See the supplementary website for videos.
@@ 4.4.3 Ablation studies
In Tab. 2 for the DAVIS I2V task, we compare our method with a variant that excludes motion conditioning using warped noise (“Ours (prompt only)”), relying solely on textual prompts describing the objects. We observe: (1) Better video reconstruction: Our method, which incorporates motion conditioning, achieves superior FID, FVD, and CoTracker mIoU scores, indicating more accurate reconstruction of the source video. This is because textual prompts and the initial frame alone are insufficient to capture a video’s future dynamics, whereas incorporating real video-derived motion guidance enables the generation of more realistic sequences. (2) Improved video quality: By utilizing warped noise for motion conditioning, our approach not only maintains but also enhances overall video quality, as measured by Vbench, demonstrating that integrating realistic motion cues improves the plausibility of the generated videos without compromising quality.
We perform additional ablation studies for the DL3DV I2V task of camera movement transfer application. Our final noise-warping-controlled model is trained using a LoRA with rank 2048. For one ablation (“Ours (lower LoRA rank)”) in Tab. 2, we fine-tune on a lower LoRA rank (256) for reduced memory consumption and improved efficiency, finding that performance deteriorates across all metrics. Especially, with FID and FVD significantly affected. This is because lighter-weight LoRA fine-tuning is unable to effectively adapt to the new warped noise distribution.
Additionally, we find in Tab. 2 that, while fine-tuning and testing without noise degradation (“Ours w/o (noise degradation)”) improves the motion controllability of driving videos (increasing CoTracker mIoU), the other metrics tend to worsen. Sometimes objects fade in and out of existence at different points in the generated videos. This is because, with a lower degradation value, the input signal of the warped noise is more strongly preserved. However, because the optical flow is not always accurate, this over-reliance on warped noise can cause quality deterioration at test time if the given flow does not match the first frame. Noise degradation during fine-tuning serves as augmentation for improved robustness, while at inference, it allows the diffusion model to draw from its priors, decreasing the strength of our motion conditioning and allowing for looser flow control. See Fig. 4 for an example.
@@ 7 Qualitative results of training-free image diffusion based video editing
Figure 11: A direct visualization of the noise produced by our noise warping algorithm, HIWYN [8], bilinear, and nearest neighbor interpolations. The forward movement in this long roller-coaster video forces the noise to expand significantly. Early in the video, the HIWYN baseline produces visibly non-Gaussian results. See the full video on our supplementary website.
@@ 11 Model Agnostic
Our method is data- and model-agnostic. It can be used to add motion control to arbitrary video diffusion models by only processing the noise sampling during fine-tuning. For example, it also works with AnimateDiff [16] fine-tuned on the WebVid dataset [2]. See its qualitative results in Fig. 16. Therefore, our method will generalize to future more advanced video diffusion base model.

### v4

@@ 4.4.1 Local object motion control
We also provide quantitative benchmarks. Following [39], we use the VIPSeg [37] to benchmark our method on local object motion control, as well as the 40 videos from our user study.
Baselines. We evaluate our video generation model against five state-of-the-art baselines, SG-I2V [39], MotionClone [34], DragAnything [60], to benchmark its ability to accurately control object and camera movements derived from a given input template. One of the most recent works, SG-I2V, is an I2V model for object motion transfer guided by bounding box trajectories. We adapt our user-defined polygons to bounding boxes as its input.
Results. From  Fig. 5, Fig. 6, Tab. 2 and our webpage, we observe:
@@
(2) We find that SG-I2V frequently misinterprets object-specific movements as global camera shifts, resulting in scene-wide translations rather than accurate object manipulations.
@@
(5) Qualitatively, our model outperforms these baselines by maintaining high object fidelity and 3D consistency, even in scenarios with intricate or overlapping motions. Notably, our approach preserves object integrity and introduces plausible physical interactions, such as generating realistic splashes when moving a duck within a tub. Extensive user studies and quantitative evaluations validate our superior performance in motion consistency, visual fidelity, and overall realism.
(6) Our quantitative evaluation matches our qualitative observations. On both VIPSeg and the 40 videos from our user-study, our method outperforms all the training-based and training-free baselines.
@@ 4.4.2 Motion transfer and camera movement control
Table 2: Quantitative comparisons of motion transfer. $\Uparrow$/$\Downarrow$ indicates a higher/lower value is better. Bold indicates the best results. Gray background rows indicate our final model. Dashed lines separate ablation study from baseline benchmarking.
CoTracker Optical Pixel CLIP CLIP FVD VBench $\Uparrow$
Training FID mIoU flow MSE text image ×$\times 10^{3}$ Subject Background Motion Temperal
free? $\Downarrow$ $\Uparrow$ err. $\Downarrow$ $\Downarrow$ $\Uparrow$ $\Uparrow$ $\Downarrow$ consistency consistency smoothness flickering
Local object motion control on VIPSeg
MotionClone 85.2 0.71 0.48 0.086 0.31 0.95 1.26 0.88 0.85 0.94 0.90
SG-I2V 61.4 0.63 0.84 0.065 0.31 0.97 1.06 0.93 0.95 0.96 0.94
Ours $\times$ 41.1 0.75 0.36 0.039 0.32 0.98 0.47 0.91 0.92 0.97 0.95
Local object motion control on our 40 samples in the user study
MotionClone 96.6 - 0.80 0.048 0.33 0.98 1.38 0.86 0.93 0.97 0.95
SG-I2V 79.9 - 0.64 0.042 0.32 0.98 1.27 0.95 0.95 0.98 0.94
DragAnything $\times$ 82.8 - 0.62 0.047 0.31 0.97 1.30 0.93 0.95 0.98 0.95
Ours $\times$ 74.3 - 0.56 0.028 0.32 0.98 0.94 0.96 0.95 0.98 0.96
Motion transfer T2V on DAVIS
DMT - 0.85 0.28 - 0.31 0.95 - 0.86 0.92 0.94 0.91
MotionClone - 0.75 0.38 - 0.32 0.93 - 0.78 0.89 0.86 0.81
MotionCtrl $\times$ 0.47 0.85 - 0.32 0.97 - 0.97 0.93 0.98 0.92
Ours $\times$ - 0.70 0.41 - 0.33 0.98 - 0.88 0.93 0.97 0.89
[…]
Motion transfer I2V on DAVIS
MotionClone 99.4 0.72 0.42 0.068 0.31 0.94 1.84 0.75 0.85 0.92 0.87
ImageConductor $\times$ 104.6 0.66 0.64 0.072 0.31 0.93 1.58 0.77 0.88 0.93 0.90
Original CogVideoX-5B 76.62 0.52 0.67 0.088 0.31 0.96 1.36 0.85 0.91 0.96 0.92
Ours ($\gamma=0.5$) $\times$ 78.6 0.74 0.36 0.053 0.31 0.97 1.21 0.88 0.92 0.98 0.93
[…]
Our ($\gamma=0.8$) $\times$ 80.6 0.68 0.47 0.067 0.31 0.96 1.50 0.85 0.91 0.96 0.92
Our ($\gamma=0.4$) $\times$ 77.7 0.74 0.36 0.056 0.31 0.97 1.27 0.87 0.91 0.97 0.93
Our ($\gamma=0.2$) $\times$ 77.1 0.74 0.37 0.058 0.32 0.97 1.29 0.86 0.91 0.97 0.93
[…]
Our (12.5% data) $\times$ 105.2 0.71 0.39 0.072 0.31 0.96 1.93 0.84 0.89 0.97 0.91
Camera movement transfer I2V on DL3DV
MotionClone 82.7 0.71 0.44 0.104 0.33 0.94 1.11 0.74 0.85 0.91 0.86
ImageConductor $\times$ 89.2 0.61 0.78 0.068 0.31 0.95 0.91 0.85 0.90 0.95 0.93
Ours $\times$ 48.4 0.83 0.20 0.046 0.32 0.97 0.34 0.88 0.92 0.97 0.93
Camera movement transfer I2V on WonderJourney
MotionClone 177.9 0.81 0.17 0.103 0.32 0.96 1.93 0.75 0.87 0.93 0.87
ImageConductor $\times$ 166.1 0.79 0.39 0.085 0.32 0.94 1.63 0.79 0.88 0.93 0.90
Ours $\times$ 128.3 0.85 0.15 0.072 0.31 0.98 1.55 0.82 0.91 0.98 0.94
Baselines. For the motion transfer T2V scenario, we compare with the recent state-of-the-art methods Diffusion Motion Transfer (DMT) [65], MotionClone [34], and MotionCtrl [58]. For the motion transfer I2V, we compare MotionClone and ImageConductor [32]) as DMT does not take an image as input.
@@
Results. We present both qualitative and quantitative comparisons with baselines in Tab. 2, Fig. 7, Fig. 8, Fig. 9, Fig. 10, and our webpage. We observe:
@@
(2) Our superior camera movement control: On the DL3DV and WonderJourney datasets, which involve substantial camera movement, our method notably outperforms MotionClone in both motion fidelity and general video quality. This highlights our method’s ability to effectively replicate intricate camera movements while maintaining visual coherence. For our depth-warping example Fig. 9, our results are far better than simply warping an image from its depth map, resulting in a smooth, realistic camera trajectory. See our webpage for videos.
@@ 4.4.3 Ablation studies
In Tab. 2 for the DAVIS I2V task, we compare our method with a variant that excludes motion conditioning using warped noise (“Original CogVideoX-5B”), relying solely on textual prompts describing the objects. We observe: (1) Better video reconstruction: Our method, which incorporates motion conditioning, achieves superior FID, FVD, and CoTracker mIoU scores, indicating more accurate reconstruction of the source video. This is because textual prompts and the initial frame alone are insufficient to capture a video’s future dynamics, whereas incorporating real video-derived motion guidance enables the generation of more realistic sequences. (2) Improved video quality: By utilizing warped noise for motion conditioning, our approach not only maintains but also enhances overall video quality, as measured by Vbench, demonstrating that integrating realistic motion cues improves the plausibility of the generated videos without compromising quality.
Further exploring the effects of degradation, we find that as the degradation values $\gamma$ increase, the motion control becomes tighter - resulting in higher optical flow and CoTracker-mIoU scores, along with a closer per-frame similarity to target videos on the I2V DAVIS task. We find that, in general, $\gamma\approx 0.5$ is a good value for most tasks.
We also perform ablations on dataset size, comparing models trained on different fractions of our dataset: training with a fraction of dataset yields worse performance than our full model.
In addition, we perform an ablation where we use the weaker base model CogVideoX-2B, and find its performance is weaker than our main T2V model, based on CogVideoX-5B.
@@ 7 Qualitative results of training-free image diffusion based video editing
Figure 11: A direct visualization of the noise produced by our noise warping algorithm, HIWYN [8], bilinear, and nearest neighbor interpolations. The forward movement in this long roller-coaster video forces the noise to expand significantly. Early in the video, the HIWYN baseline produces visibly non-Gaussian results. See the full video on our webpage.
@@ 11 Model Agnostic
Our method is data- and model-agnostic. It can be used to add motion control to arbitrary video diffusion models by only processing the noise sampling during fine-tuning. For example, it also works with AnimateDiff [16] fine-tuned on the WebVid dataset [2] (the weights for this model on our GitHub page). See its qualitative results in Fig. 16. Since release, the community has also trained a version of Go-with-the-Flow on HunyuanVideo (linked on our GitHub page). Therefore, our method will generalize to future more advanced video diffusion base model.